300 gms of sugar solution contains 40% sugar. How much sugar should be added to make it 50% sugar solution?
Answer by Earlsdon(6294)

Let x = the number of grams of sugar to be added to the 300gm of 40% sugar solution.
300(0.4)+x = (300+x)(0.5) Simplify.
120+x = 150+0.5x Subtract 0.5x from both sides.
120+0.5x = 150 Now subtract 120 from both sides.
0.5x = 30 Multiply both sides by 2.
x = 60
Add 60 grams of sugar to the 300gm of 40% sugar solution to obtain 360gm of 50% sugar solution.